# Comparative Performance Analysis of Crystals in Total-Body PET Scanners: Monte-Carlo Simulation Study with Different Materials and Geometry

### v1

@@ 1.2 Scintillator Configurations
For all the available crystals, LYSO has gained popularity for its high density, good light yield, and faster scintillation decay time [15, 16]. There are still concerns regarding its intrinsic radioactivity, high cost, and potential time and energy resolution improvement. Therefore, ongoing research explores alternative materials and geometries to enhance scanner performance [17, 6]. Among these, the multi-component garnet $Gd_{3}Al_{2}Ga_{3}O_{12}$ doped with Cerium and Magnesium (GAGG) is one such promising material [18]. While Mg co-doping does not enhance its light yield, it expedites the scintillation time response of the scintillator, which is crucial in TOF-related studies of the PET scanner [19, 20, 21]. We have used Mg,Ce:GAGG for its high light yield, faster decay decay time, and absence of internal radioactivity, which could make it a better alternative to LYSO. Properties of both the materials are given in table 1. The detector crystal’s shape also influences the detector’s timing property. The scintillating photons reflected off the pyramid surface will alter their trajectory and directly reach the coupling surface, enhancing light output and improving timing resolution. Cuboid-shaped crystals often require multiple photon reflections for absorption, potentially impacting timing resolution. To mitigate this, we explored pyramid-shaped crystals [22]. The simulated images of both of the crystal shapes is shown in figure 1.
@@ 2.1 Simulation Setup
For the sensitivity test, we will be constructing three PET scanners with geometrical parameters of uEXPLORER. Those are a) Actual uEXPLORER consisting LYSO crystals of length 18mm. b) 0.1% Mg, 1% Ce Doped GAGG crystals of length 18mm. c) Pyramid-shaped LYSO crystal of length 18-mm. The geometrical parameters of the PET scanner are given in table 2, and its visual representation is given in figure 2. The LYSO crystal has a density of 7.11 g/cm3, and its composition includes 71.447% of Lutetium, 4.034% of Yttrium, 6.371% of Silicon, and 18.148% of Oxygen [12]. The pyramid crystals were simulated for 18-mm length. This geometry of detector is a trade-off between the CTR and sensitivity.
Table 1: Properties of LSO and Mg,Ce:GAGG crystals [26, 20]
Properties LSO 0.1% Mg, 1% Ce doped GAGG
@@
The Geant4 simulation toolkit was employed to mimic the actual performance of the scanner. Physics list G4EmStandardPhysics_option3, G4DecayPhysics, and G4RadioactiveDecayPhysics were used to simulate the underlying physics involved. The optical physics module was disabled for this setup to save memory and system load. Then, a blurring module was used which assigned energy resolution to the simulated crystals. A mean energy resolution value of 11.3% for LYSO and 8.6% in GAGG [27, 21] in reference to 511 keV was adopted, whereas a detection efficiency factor of 0.883 was added to the recorded pulses. Next, we conducted NEMA performance tests, the standard method for assessing PET system performance and comparing it across different systems. In this work, the NEMA-NU2-2018 standards [28] are used to first validate the sensitivity for uEXPLORER and subsequently evaluate the performance of other scanners on the same geometry.
@@ 3.2 Intrinsic Coincident Time Resolution
The Full Width at Half Maximum (FWHM) of the time difference distribution of the first photon hits represents the intrinsic Coincidence Time Resolution (CTR) of our system. This intrinsic CTR is a core measure of the detector’s temporal performance, as it remains unaffected by external factors like electronic noise in the readout circuits, timing jitter in the photodetectors, variations in the crystal growth and surface treatment, temperature fluctuations, etc., Therefore, the CTR value we obtain is expected to be lower than what would be observed in real-world experimental conditions, as shown in figure 6. The Mg,Ce:GAGG cuboidal crystal has an intrinsic CTR of 114 ps, compared to 107 ps for the LYSO cuboidal crystal. For the pyramid-shaped crystals, Mg,Ce:GAGG shows a very slight deterioration in CTR by 0.3 ps compared to LYS;O. As this difference is minimal, the pyramid Mg,CeGAGG data was not shown separately in the plot. Overall, pyramid-shaped crystals exhibit a 60.6% improvement in CTR as compared to cuboidal crystals.
@@
3.3 Energy Resolution
In figure 7, the x-axis represents the number of photons hitting the sensitive detector in each event. Mg,Ce:GAGG showed an energy resolution of 8.59% for cuboidal crystals and 7.21% for pyramid-shaped crystals, reflecting a 16.06% improvement due to the change in shape. Similarly, LYSO had an energy resolution of 11.57% for cuboidal crystals and 10.33% for pyramid-shaped crystals, indicating a 10.3% improvement with the shape modification. The better performance of pyramid crystals is due to better Light Collection Efficiency (LCE) in this configuration. This shape minimizes optical-photon reflections at the scintillator-reflector boundary, thereby increasing the likelihood of photon detection at the photocathode and enhancing energy resolution [7]. Overall, Mg,Ce:GAGG demonstrates much better energy resolution than the LYSO crystal.
@@ 4 Conclusion
In this work, we demonstrate that pyramid-shaped crystals of LYSO can improve intrinsic CTR by ~60% and energy resolution by ~11.6% compared to traditional cuboidal crystals, indicating their potential usage in TOF PET. Mg,Ce:GAGG demonstrated CTR of 114 ps, which is rather good (though not better) if compared with 107 ps of LYSO, outperforming many high-density scintillators like BGO. Mg,Ce:GAGG-cuboid also shows 25.7% better energy resolution than LYSO-cuboid. Thus, The Mg, Ce-doped GAGG crystal emerges as a promising detector material due to its high light yield, finer energy resolution, and lack of internal radioactivity. Although both pyramid-shaped crystals and Mg,Ce:GAGG crystals exhibit reduced sensitivity, this limitation can be addressed using depth of interaction (DOI) techniques. In conclusion, both pyramid-shaped crystals and Mg, Ce doped GAGG crystals show great promise for PET applications and open the way for optimizing crystal configurations within PET scanners towards better image quality and quantitative accuracy in research.

### v2

@@ 1.2 Scintillator Configurations
For all the available crystals, LYSO has gained popularity for its high density, good light yield, and faster scintillation decay time [15, 16]. There are still concerns regarding its intrinsic radioactivity, high cost, and potential time and energy resolution improvement. Therefore, ongoing research explores alternative materials and geometries to enhance scanner performance [17, 6]. Among these, the multi-component garnet $Gd_{3}Al_{2}Ga_{3}O_{12}$ doped with Cerium and Magnesium (GAGG) is one such promising material [18]. While Mg co-doping does not enhance its light yield, it expedites the scintillation time response of the scintillator, which is crucial in TOF-related studies of the PET scanner [19, 20, 21]. We have used Mg,Ce:GAGG for its high light yield, faster decay decay time, and absence of internal radioactivity, which could make it a better alternative to LYSO. Moreover, Mg,Ce:GAGG also shows superior radiation hardness as compared to LYSO [22]. Properties of both the materials are given in table 1. The detector crystal’s shape also influences the detector’s timing property. The scintillating photons reflected off the pyramid surface will alter their trajectory and directly reach the coupling surface, enhancing light output and improving timing resolution. Cuboid-shaped crystals often require multiple photon reflections for absorption, potentially impacting timing resolution. To mitigate this, we explored pyramid-shaped crystals [23]. While pyramid shapes can enhance light collection, the variations in the crystal’s cross-sectional area along its height can introduce position-dependent signal variations, which are needed to be considered. However detailed characterization of this effect in pyramid-shaped crystals is still limited in the literature. The simulated images of both of the crystal shapes is shown in figure 1.
@@ 2.1 Simulation Setup
For the sensitivity test, we will be constructing three PET scanners with geometrical parameters of uEXPLORER. Those are a) Actual uEXPLORER consisting LYSO crystals of length 18mm. b) 0.1% Mg, 1% Ce Doped GAGG crystals of length 18mm. c) Pyramid-shaped LYSO crystal of length 18-mm. The pyramidal geometry of the detector is a trade-off between the CTR and sensitivity. With increasing crystal length, sensitivity improves but CTR deteriorates [27, 28]; therefore, 18-20 mm crystal length is widely used for a better tradeoff between these two parameters. We study this correlation in pyramidal crystals as shown in figure 2. The geometrical parameters of the PET scanner are given in table 2, and its visual representation is given in figure 3. The LYSO crystal has a density of 7.11 g/cm3, and its composition includes 71.447% of Lutetium, 4.034% of Yttrium, 6.371% of Silicon, and 18.148% of Oxygen [12]. The pyramid crystals were simulated for 18-mm length.
Table 1: Properties of LYSO and Mg,Ce:GAGG crystals [29, 20]
Properties LYSO 0.1% Mg, 1% Ce doped GAGG
@@
Figure 2: Sensitivity and intrinsic CTR values for pyramidal crystals of different lengths.
@@
The Geant4 simulation toolkit was employed to mimic the actual performance of the scanner. Physics list G4EmStandardPhysics_option3, G4DecayPhysics, and G4RadioactiveDecayPhysics were used to simulate the underlying physics involved. The optical physics module was disabled for this setup to save memory and system load. Then, a blurring module was used which assigned energy resolution to the simulated crystals. This module introduced Gaussian energy blur to the absorbed energy. It allowed us to achieve an energy resolution matching the experimental energy resolution. So, mean energy resolution value of 11.7% [12] for LYSO and 8.6% in GAGG [30, 21] in reference to 511 keV was adopted, whereas a detection efficiency factor of 0.883 was added to the recorded pulses. Next, we conducted NEMA performance tests, the standard method for assessing PET system performance and comparing it across different systems. In this work, the NEMA-NU2-2018 standards [31] are used to first validate the sensitivity for uEXPLORER and subsequently evaluate the performance of other scanners on the same geometry.
@@ 3.2 Intrinsic Coincident Time Resolution
The Full Width at Half Maximum (FWHM) of the time difference distribution of the first photon hits represents the intrinsic Coincidence Time Resolution (CTR) of our system. This intrinsic CTR is a core measure of the detector’s temporal performance, as it remains unaffected by external factors like electronic noise in the readout circuits, timing jitter in the photodetectors, variations in the crystal growth and surface treatment, temperature fluctuations, etc., Therefore, the CTR value we obtain is expected to be lower than what would be observed in real-world experimental conditions, as shown in figure 7. The Mg,Ce:GAGG cuboidal crystal has an intrinsic CTR of 114 ps, compared to 107 ps for the LYSO cuboidal crystal. For the pyramid-shaped crystals, Mg,Ce:GAGG shows a very slight deterioration in CTR by 0.3 ps compared to LYSO. As this difference is minimal, the pyramid Mg,CeGAGG data was not shown separately in the plot. Overall, pyramid-shaped crystals exhibit a 60.6% improvement in CTR as compared to cuboidal crystals.
@@
3.3 Energy Resolution
In figure 8, the x-axis represents the number of photons hitting the sensitive detector in each event. Mg,Ce:GAGG showed an energy resolution of 8.59% for cuboidal crystals and 7.21% for pyramid-shaped crystals, reflecting a 16.06% improvement due to the change in shape. Similarly, LYSO had an energy resolution of 11.57% for cuboidal crystals and 10.33% for pyramid-shaped crystals, indicating a 10.71% improvement with the shape modification. The better performance of pyramid crystals is due to better Light Collection Efficiency (LCE) in this configuration. This shape minimizes optical-photon reflections at the scintillator-reflector boundary, thereby increasing the likelihood of photon detection at the photocathode and enhancing energy resolution [7]. Overall, Mg,Ce:GAGG demonstrates much better energy resolution than the LYSO crystal.
@@ 4 Conclusion
In this work, we demonstrate that pyramid-shaped crystals of LYSO can improve intrinsic CTR by ~60.6% and energy resolution by ~10.71% compared to traditional cuboidal crystals, indicating their potential usage in TOF PET. Mg,Ce:GAGG demonstrated CTR of 114 ps, which is rather good (though not better) if compared with 107 ps of LYSO, outperforming many high-density scintillators like BGO. Mg,Ce:GAGG-cuboid also shows 25.7% better energy resolution than LYSO-cuboid. Thus, The Mg,Ce-doped GAGG crystal emerges as a promising detector material due to its high light yield, finer energy resolution, and lack of internal radioactivity. Although both pyramid-shaped crystals and Mg,Ce:GAGG crystals exhibit reduced sensitivity, this limitation can be addressed using depth of interaction (DOI) techniques. In conclusion, both pyramid-shaped crystals and Mg, Ce doped GAGG crystals show great promise for PET applications and open the way for optimizing crystal configurations within PET scanners towards better image quality and quantitative accuracy in research.